# Geometric foundations of thermodynamics in the quantum regime

### v1

@@ VII Distances
In quantum theory, the Bures-Wasserstein distance defines a distance measure comparing quantum states, represented by density matrices, by considering the properties of positive-definite self-adjoint matrices [oostrum22_bures]. This distance is particularly suited for the quantum thermodynamic state space, as it aligns with the Riemannian geometry of the space of quantum states $\mathcal{D}$.
@@
The above definition can be shown to be independent of the specific curve chosen [caja12_variedades].
@@
The proof is detailed in [oostrum22_bures]. In brief, the metric $g_{\mathrm{BW}}$ arises from the infinitesimal form of the Bures-Wasserstein distance, where $L_{X}$ solves the Lyapunov equation for the perturbation $X(\rho_{\boldsymbol{\lambda}})$. The real part ensures symmetry, and the trace preserves positive-definiteness, establishing $g_{\mathrm{BW}}$ as a Riemannian metric. ∎
@@
$g_{ij}(\boldsymbol{\lambda})=\textrm{Re}\left[\operatorname{\mathrm{tr}}\left(\rho_{\boldsymbol{\lambda}}L_{i}L_{j}\right)\right],\quad L_{i}=\frac{\partial\ln(\rho_{\boldsymbol{\lambda}})}{\partial\lambda_{i}},$ (34)
with the line element
@@ VIII Geodesics and quasistatic processes
$\Sigma_{\gamma}(t)=\int_{0}^{T}\varsigma(t)dt=\int_{0}^{T}\kappa g_{\mathrm{BW}}(\gamma^{\prime}(t),\gamma^{\prime}(t))dt.$ (38)
For finite-speed processes, $\varsigma_{\gamma}(t)>0$ due to non-zero velocity $\gamma^{\prime}(t)$, indicating irreversibility [tejero2025entropy]. In the quasistatic limit ($\lambda_{i}^{\prime}\to 0$), $\varsigma_{\gamma}(t)\to 0$ and $\Sigma_{\gamma}(t)\to 0$. Minimizing geodesics minimize both $L(\gamma)$ and $\Sigma$, optimizing the path to reduce the entropy production, thus achieving maximal reversibility.
Remark 12.
The term $g_{\mathrm{BW}}(\gamma^{\prime},\gamma^{\prime})$ measures the instantaneous rate of state evolution, driving irreversibility via excitations or non-equilibrium effects. Minimizing geodesics mitigate these by following the smoothest path in state space, allowing the system to remain in equilibrium at each step, analogous to adiabatic transformations. The proportionality $\varsigma_{\gamma}\propto g_{\mathrm{BW}}(\gamma^{\prime},\gamma^{\prime})$ reflects that faster transformations disrupt equilibrium more severely, increasing $\Sigma_{\gamma}$.
The following results formalize this connection between geodesics and quasistatic processes.
@@
Quasistatic processes evolve through Gibbs states $\rho_{\boldsymbol{\lambda}}$ with minimal dissipation. The Bures-Wasserstein metric quantifies state distinguishability, and minimizing geodesics minimize thermodynamic length $L(\gamma)$, corresponding to paths of least entropy production $\Sigma_{\gamma}$, as required for optimal quasistatic transformations. ∎
@@
This result is a direct consequence of the Hopf-Rinow theorem. Since $(\mathcal{B},g_{\mathrm{BW}})$ is a Riemannian manifold, by the Hopf-Rinow theorem completeness of $(\mathcal{B},g_{\mathrm{BW}})$ implies that any two points can be joined by a minimizing geodesic. ∎
@@ IX Boundary effects
    For any smooth curve $\gamma:[0,1)\to\mathcal{D}^{\circ}$ such that $\lim_{t\to 1^{-}}\gamma(t)=\overline{\rho}\in\partial\mathcal{D}$ with $\mathrm{rank}(\overline{\rho})=k<m$, the entropy satisfies
    $\lim_{t\to 1^{-}}S(\gamma(t))=S(\overline{\rho}),$ (39)
    where $S(\overline{\rho})=-\sum_{i=1}^{k}\overline{p}_{i}\ln\overline{p}_{i}$, and $\overline{p}_{i}$ with $i=1,\dots,k$ (not necessarily all distinct) are the positive eigenvalues of $\overline{\rho}$, where $k$ is the rank of $\overline{\rho}$. In particular, if $\overline{\rho}$ is a pure state $(k=1)$, then $S(\gamma(t))\to 0$, and if $\overline{\rho}$ is maximally mixed on its support, i.e. $\overline{p}_{i}=1/k$ for $i=1,\ldots,k$, then $S(\gamma(t))\to\ln k$.
@@
First, we establish the continuous extension of $S$ to $\partial\mathcal{D}$. For a state $\overline{\rho}\in\mathcal{D}$ with eigenvalues $p_{i}\geq 0$, some possibly zero, the entropy is defined by extending the function $f(x)=-x\ln x$ to the domain boundary $x=0$. Let $\gamma:[0,1)\to\mathcal{D}^{\circ}$ be a smooth curve with $\lim_{t\to 1^{-}}\gamma(t)=\overline{\rho}\in\partial\mathcal{D}$ of rank $k<m$. In a neighborhood of $\overline{\rho}$, $\gamma(t)$ can be diagonalized: $\gamma(t)=U(t)\operatorname{diag}(p_{1}(t),\ldots,p_{m}(t))U^{*}(t)$ with $p_{i}(t)>0$, $\sum_{i=1}^{m}p_{i}(t)=1$, and $U(t)$ is a $m\times m$ unitary matrix for each value of $t$. As $t\to 1^{-}$, assume without loss of generality that $p_{i}(t)\to\overline{p}_{i}>0$ for $i=1,\ldots,k$, and $p_{i}(t)\to 0$ for $i=k+1,\ldots,m$, with $\sum_{i=1}^{k}\overline{p}_{i}=1$. Then
@@
The first sum converges to $S(\overline{\rho})$ by continuity of $f(x)$ at $\overline{p}_{i}>0$. For the second sum, consider $f(p_{i}(t))=-p_{i}(t)\ln p_{i}(t)$. As $p_{i}(t)\to 0^{+}$, $f(p_{i}(t))\to 0$. The sum $\sum_{i=k+1}^{m}f(p_{i}(t))$ has mkm-k terms and each term converges to 0, so $\sum_{i=k+1}^{m}p_{i}(t)\ln p_{i}(t)\to 0$. Thus,
$\lim_{t\to 1^{-}}S(\gamma(t))=S(\overline{\rho}).$ (41)
The special cases follow immediately: pure states give $S(\overline{\rho})=0$; maximal mixing on support gives $S(\overline{\rho})=\ln k$. ∎
@@
where $k$ is the rank of the target boundary stratum. Moreover, boundary strata, are unattainable along any finite-length geodesic in $\mathcal{B}$.
@@ X Connections, curvature and holonomy in the quantum thermodynamic fiber bundle
The fiber $F_{\sigma}=\Xi^{-1}(\sigma)$ contains all thermodynamic configurations yielding the fixed quantum state $\sigma\in\mathcal{D}$, with only the equilibrium point $p_{\sigma}=F_{\sigma}\cap\mathcal{E}$ lying on the Legendrian submanifold $\mathcal{E}$. Geodesics on $\mathcal{E}$ connect equilibrium states across different fibers, while non-equilibrium points in $F_{\sigma}\setminus\mathcal{E}$ represent transient configurations. Relaxation paths within $F_{\sigma}$ converge to $p_{\sigma}$, and their dissipation can be quantified and minimized using a pseudo-Riemannian metric $g_{\mathcal{M}}$ on $\mathcal{M}$ that extends the Bures-Wasserstein metric $g_{\mathrm{BW}}$ from $\mathcal{E}$.
Thermodynamic processes, either quasistatic evolutions along Gibbs states or driven non-equilibrium dynamics, correspond to paths in $\mathcal{M}$ projecting via $\Xi$ to curves in $\mathcal{D}$. To model these transitions geometrically, we introduce an Ehresmann connection on the fiber bundle, decomposing $T\mathcal{M}$ into vertical and horizontal subbundles. This connection enables parallel transport of thermodynamic states along paths in $\mathcal{D}$, with curvature quantifying non-integrability and holonomy inducing geometric irreversibility in cyclic processes. This is analogous to holonomies existent in gauge theories.
@@
$\mathscr{V}_{p}=\ker\left(d\Xi_{p}:T_{p}\mathcal{M}\to T_{\Xi(p)}\mathcal{D}\right)=T_{p}F_{\Xi(p)},$ (43)
the tangent space to the fiber $F_{\sigma}$ at $p\in F_{\sigma}$. Since $\dim\mathcal{M}=2n+1$ and $\dim\mathcal{D}=n$, we have $\dim\mathscr{V}_{p}=n+1$. Assuming that $\Xi$ is a submersion, $d\Xi_{p}$ is surjective, and $\mathscr{V}=\ker(d\Xi)$ is a smooth subbundle.
@@
    $d\Xi_{p}|_{\mathscr{H}_{p}}:\mathscr{H}_{p}\to T_{\Xi(p)}\mathcal{D}$ is a linear isomorphism.
@@
Let $\pi:\mathcal{M}\to\mathcal{E}$ be the projection along fibers to the equilibrium submanifold. At any point $p=(S,\boldsymbol{a},\boldsymbol{\lambda})\in\mathcal{M}$, the tangent space is spanned by
$T_{p}\mathcal{M}=\operatorname{span}\left\{\frac{\partial}{\partial S},\ \frac{\partial}{\partial a_{1}},\ \dots,\ \frac{\partial}{\partial a_{n}},\ \frac{\partial}{\partial\lambda_{1}},\ \dots,\ \frac{\partial}{\partial\lambda_{n}}\right\}.$ (44)
For $p\in F_{\sigma}$ with $\mu_{i}=\lambda_{i}$, we have $\lambda_{i}=c_{i}$, for $i=1,\ldots,n$, so
$\mathscr{V}_{p}=T_{p}F_{\sigma}=\operatorname{span}\left\{\frac{\partial}{\partial S},\ \frac{\partial}{\partial a_{1}},\ \dots,\ \frac{\partial}{\partial a_{n}}\right\},$ (45)
and the subbundle
$\mathscr{H}_{p}=\operatorname{span}\left\{\frac{\partial}{\partial\lambda_{1}},\ \dots,\ \frac{\partial}{\partial\lambda_{n}}\right\}$ (46)
projects isomorphically to $T_{p_{\sigma}}\mathcal{E}$ via $\pi_{*}$.
@@
Consider a pseudo-Riemannian metric $g_{\mathcal{M}}$ on $\mathcal{M}$ such that $g_{\mathcal{M}}|_{\mathscr{V}_{p}}$ is positive-definite. This metric induces an Ehresmann connection by defining
@@
where $\beta=1,\ldots,n+1$, for a basis $\{w_{\beta}\}$ of $\mathscr{V}_{p}$. The matrix $g_{\mathcal{M}}(w_{\alpha},w_{\beta})$ is invertible since $g_{\mathcal{M}}|_{\mathscr{V}_{p}}$ is positive-definite, ensuring a unique $v$.
@@
$d\Xi_{p}(T_{p}\mathcal{M})=d\Xi_{p}(\mathscr{H}_{p}).$ (49)
Given $\dim\mathscr{H}_{p}=\dim T_{\Xi(p)}\mathcal{D}=n$ and $\Xi$ a submersion, $d\Xi_{p}|_{\mathscr{H}_{p}}$ is an isomorphism. The smoothness of $\mathscr{H}$ follows from the smoothness of $g_{\mathcal{M}}$ and $\Xi$. ∎
Remark 14.
@@
A pseudo-Riemannian metric on $\mathcal{M}$, $g_{\mathcal{M}}$, satisfying the desiderata in Remark 14 is
$g_{\mathcal{M}}=g_{S}dS^{2}+\sum_{i=1}^{n}g_{a_{i}}da_{i}^{2}+\pi^{*}g_{\mathrm{BW}}+\sum_{i=1}^{n}h_{i}(dS\otimes d\lambda_{i}+d\lambda_{i}\otimes dS),$ (50)
where:
  • $g_{S}(p_{\sigma})\in\mathbb{R}$ controls the entropy direction, which may be negative,
  • $g_{a_{i}}(p_{\sigma})>0$ are positive-definite on expectation value directions, for all $i=1,\ldots,n$,
  • $\pi^{*}g_{\mathrm{BW}}=\sum_{i,j}g_{ij}(\pi(p_{\sigma}))d\lambda_{i}d\lambda_{j}$ pulls back the Bures-Wasserstein metric,
  • $h_{i}(p)$ are cross-terms ensuring contact compatibility, for all $i=1,\ldots,n$.
Note that, on $\mathcal{E}$, $\eta=0$ implies $dS=\sum\lambda_{i}da_{i}$, so
$g_{\mathcal{M}}|_{T\mathcal{E}}=\pi^{*}g_{\mathrm{BW}}.$ (51)
On a fiber $F_{\sigma}$, where $\lambda_{i}=c_{i}$ fixed, $d\lambda_{i}=0$ for all $i=1,\ldots,n$, and
$g_{\mathcal{M}}|_{F_{\sigma}}=g_{S}dS^{2}+\sum_{i=1}^{n}g_{a_{i}}da_{i}^{2}.$ (52)
For a path $\gamma:[0,T]\to\mathcal{M}$ with $\gamma(t)=(S(t),\boldsymbol{a}(t),\boldsymbol{\lambda}(t))$, the thermodynamic length is
$L(\gamma)=\int_{0}^{T}\sqrt{g_{\mathcal{M}}(\gamma^{\prime}(t),\gamma^{\prime}(t))}dt,$ (53)
@@
$g_{\mathcal{M}}(\gamma^{\prime},\gamma^{\prime})=g_{S}S^{\prime 2}+\sum_{i=1}^{n}g_{a_{i}}(a_{i}^{\prime})^{2}+\sum_{i,j=1}^{n}g_{ij}\left(\pi(\gamma(t))\right)\lambda_{i}^{\prime}\lambda_{j}^{\prime}+2\sum_{i=1}^{n}h_{i}S^{\prime}\lambda_{i}^{\prime}.$ (54)
Along a relaxation path $\gamma(t)\in F_{\sigma}$, with $\lambda_{i}^{\prime}=0$, for all $i=1,\ldots,n$,
$L(\gamma)=\int_{0}^{T}\sqrt{g_{S}S^{\prime 2}+\sum_{i=1}^{n}g_{a_{i}}(a_{i}^{\prime})^{2}}dt.$ (55)
@@
To ensure the compatibility with the contact structure, we need the metric $g_{\mathcal{M}}$ to be positive-definite on $\ker(\eta)$, and possibly indefinite in the transverse direction (e.g., along $\partial S$). The Reeb vector field $\mathcal{R}$ is $\mathcal{R}=\partial_{S}$, since $\eta(\mathcal{R})=1$ and $d\eta=-\sum_{i=1}^{n}d\lambda_{i}\wedge da_{i}$. The metric can assign $g_{\mathcal{M}}(\mathcal{R},\mathcal{R})=g_{S}$, with cross-terms $h_{i}$ ensuring that $g_{\mathcal{M}}$ respects the structure of the contact distribution via $d\eta$.
@@
For a smooth path $\gamma:[0,1]\to\mathcal{D}$, with $\gamma(0)=\sigma_{1}$, $\gamma(1)=\sigma_{2}$, and $p_{0}\in\Xi^{-1}(\sigma_{1})$, the horizontal lift is a curve $\tilde{\gamma}:[0,1]\to\mathcal{M}$ satisfying

### v2

@@ VII Distances
In quantum theory, the Bures-Wasserstein distance defines a distance measure comparing quantum states, represented by density matrices, by considering the properties of positive-definite self-adjoint matrices [36]. This distance is particularly suited for the quantum thermodynamic state space, as it aligns with the Riemannian geometry of the space of quantum states.
@@
The above definition can be shown to be independent of the specific curve chosen [42].
@@
The proof is detailed in [36]. In brief, the metric $g_{\mathrm{BW}}$ arises from the infinitesimal form of the Bures-Wasserstein distance, where $L_{X}$ solves the Lyapunov equation for the perturbation $X(\rho_{\boldsymbol{\lambda}})$. The real part ensures symmetry, and the trace preserves positive-definiteness, establishing $g_{\mathrm{BW}}$ as a Riemannian metric. ∎
@@
$g_{ij}(\boldsymbol{\lambda})=\textrm{Re}\left[\operatorname{\mathrm{tr}}\left(\rho_{\boldsymbol{\lambda}}L_{i}L_{j}\right)\right],\ L_{i}=\frac{\partial\ln(\rho_{\boldsymbol{\lambda}})}{\partial\lambda_{i}},$ (35)
with line element
@@ VIII Geodesics and quasistatic processes
$\Sigma(t)=\int_{0}^{T}\varsigma(t)dt=\int_{0}^{T}\kappa g_{\mathrm{BW}}(\gamma^{\prime}(t),\gamma^{\prime}(t))dt.$ (39)
For finite-speed processes, $\varsigma_{\gamma}(t)>0$ due to non-zero velocity $\gamma^{\prime}(t)$, indicating irreversibility [48]. In the quasistatic limit ($\lambda_{i}^{\prime}\to 0$), $\varsigma_{\gamma}(t)\to 0$ and $\Sigma(t)\to 0$. Minimizing geodesics minimize both $L(\gamma)$ and $\Sigma$, optimizing the path to reduce the entropy production, thus achieving maximal reversibility.
The term $g_{\mathrm{BW}}(\gamma^{\prime},\gamma^{\prime})$ measures the instantaneous rate of state evolution, driving irreversibility via excitations or non-equilibrium effects. Minimizing geodesics mitigate these by following the smoothest path in state space, thereby allowing the system to remain in equilibrium at each step, analogous to adiabatic transformations. The proportionality $\varsigma_{\gamma}\propto g_{\mathrm{BW}}(\gamma^{\prime},\gamma^{\prime})$ reflects that faster transformations disrupt equilibrium more severely, increasing $\Sigma$. The following results formalize this connection between geodesics and quasistatic processes.
@@
Quasistatic processes evolve through Gibbs states $\rho_{\boldsymbol{\lambda}}$ with minimal dissipation. The Bures-Wasserstein metric quantifies state distinguishability, and minimizing geodesics minimize thermodynamic length $L(\gamma)$, corresponding to paths of least entropy production $\Sigma$, as required for optimal quasistatic transformations. ∎
@@
This result is a direct consequence of the Hopf-Rinow theorem. Since $(\mathcal{B},g_{\mathrm{BW}})$ is a Riemannian manifold, by the Hopf-Rinow theorem, completeness of $(\mathcal{B},g_{\mathrm{BW}})$ implies that any two points can be joined by a minimizing geodesic. ∎
@@ IX Boundary effects
In this section, we analyze the consequences of the boundary effects in the manifold of quantum states $\mathcal{D}$. While the core construction uses the finite-dimensional manifold $\mathcal{B}$ of Gibbs states generated by the chosen observables, the framework extends naturally to the full interior $\mathcal{D}^{\circ}$. Only in this section, we adopt $\mathcal{D}$ instead of the base $\mathcal{B}$ in order to account for the difference between the interior of the manifold and its boundary.
@@
    For any smooth curve $\gamma:[0,1)\to\mathcal{D}^{\circ}$ such that $\lim_{t\to 1^{-}}\gamma(t)=\bar{\rho}\in\partial\mathcal{D}$ with $\mathrm{rank}(\bar{\rho})=k<m$, the entropy satisfies
    $\lim_{t\to 1^{-}}S(\gamma(t))=S(\bar{\rho}),$ (40)
    where $S(\bar{\rho})=-\sum_{i=1}^{k}\bar{p}_{i}\ln\bar{p}_{i}$, and $\bar{p}_{i}$ with $i=1,\dots,k$ (not necessarily all distinct) are the positive eigenvalues of $\bar{\rho}$, where $k$ is the rank of $\bar{\rho}$. In particular, if $\bar{\rho}$ is a pure state $(k=1)$, then $S(\gamma(t))\to 0$, and if $\bar{\rho}$ is maximally mixed on its support, i.e., $\bar{p}_{i}=1/k$ for $i=1,\ldots,k$, then $S(\gamma(t))\to\ln k$.
@@
First, we establish the continuous extension of $S$ to $\partial\mathcal{D}$. For a state $\bar{\rho}\in\mathcal{D}$ with eigenvalues $p_{i}\geq 0$, some possibly zero, the entropy is defined by extending the function $f(x)=-x\ln x$ to the domain boundary $x=0$. Let $\gamma:[0,1)\to\mathcal{D}^{\circ}$ be a smooth curve with $\lim_{t\to 1^{-}}\gamma(t)=\bar{\rho}\in\partial\mathcal{D}$ of rank $k<m$. In a neighborhood of $\bar{\rho}$, $\gamma(t)$ can be diagonalized: $\gamma(t)=U(t)\operatorname{diag}(p_{1}(t),\ldots,p_{m}(t))U^{*}(t)$ with $p_{i}(t)>0$, $\sum_{i=1}^{m}p_{i}(t)=1$, and $U(t)$ is a $m\times m$ unitary matrix for each value of $t$. As $t\to 1^{-}$, assume without loss of generality that $p_{i}(t)\to\bar{p}_{i}>0$ for $i=1,\ldots,k$, and $p_{i}(t)\to 0$ for $i=k+1,\ldots,m$, with $\sum_{i=1}^{k}\bar{p}_{i}=1$. Then
@@
The first sum converges to $S(\bar{\rho})$ by continuity of $f(x)$ at $\bar{p}_{i}>0$. For the second sum, consider $f(p_{i}(t))=-p_{i}(t)\ln p_{i}(t)$. As $p_{i}(t)\to 0^{+}$, $f(p_{i}(t))\to 0$. The sum $\sum_{i=k+1}^{m}f(p_{i}(t))$ has mkm-k terms and each term converges to 0, so $\sum_{i=k+1}^{m}p_{i}(t)\ln p_{i}(t)\to 0$. Thus,
$\lim_{t\to 1^{-}}S(\gamma(t))=S(\bar{\rho}).$ (42)
The special cases follow immediately: pure states give $S(\bar{\rho})=0$; maximal mixing on support gives $S(\bar{\rho})=\ln k$. ∎
@@
where $k$ is the rank of the target boundary stratum. Moreover, boundary strata are unattainable along any finite-length geodesic in $\mathcal{B}$.
@@
The continuity of entropy to the boundary and the geometric unattainability of rank-deficient states provide a differential-geometric derivation of the third law in quantum thermodynamics. The divergence of geodesic lengths toward the boundary implies an infinite thermodynamic cost to reach low-entropy states. Consequences include fundamental limits on finite-time thermodynamic processes, where complete purification, i.e., transforming a thermal state into a pure state, is impossible thermodynamically.
@@ X Connections, curvature and holonomy in the quantum thermodynamic fiber bundle
X.1 Ehresmann connection
The fiber $F_{\sigma}=\Xi^{-1}(\sigma)$ contains all thermodynamic configurations for the fixed quantum state $\sigma\in\mathcal{B}$, with only the equilibrium point $p_{\sigma}=F_{\sigma}\cap\mathcal{B}$ lying on the Legendrian submanifold $\mathcal{E}$. Geodesics on $\mathcal{E}$ connect equilibrium states across different fibers, while non-equilibrium points in $F_{\sigma}\setminus\mathcal{E}$ represent transient configurations. Relaxation paths within $F_{\sigma}$ converge to $p_{\sigma}$, and their dissipation can be quantified and minimized using a pseudo-Riemannian metric $g_{\mathcal{M}}$ on $\mathcal{M}$ that extends the Bures-Wasserstein metric $g_{\mathrm{BW}}$ from $\mathcal{E}$.
Thermodynamic processes, either quasistatic evolutions along Gibbs states or driven non-equilibrium dynamics, correspond to paths in $\mathcal{M}$ projecting via $\Xi$ to curves in $\mathcal{B}$. To model these transitions geometrically, we introduce an Ehresmann connection on the fiber bundle, decomposing $T\mathcal{M}$ into vertical and horizontal subbundles. This connection enables parallel transport of thermodynamic states along paths in $\mathcal{B}$, with curvature quantifying non-integrability and holonomy inducing geometric irreversibility in cyclic processes. This is analogous to the holonomies existent in gauge theories.
@@ X.1 Ehresmann connection
$\mathscr{V}_{p}=\ker\left(d\Xi_{p}:T_{p}\mathcal{M}\to T_{\Xi(p)}\mathcal{B}\right)=T_{p}F_{\Xi(p)},$ (44)
the tangent space to the fiber $F_{\sigma}$ at $p\in F_{\sigma}$. Since $\dim\mathcal{M}=2n+1$ and $\dim\mathcal{B}=n$, we have $\dim\mathscr{V}_{p}=n+1$. Assuming that $\Xi$ is a submersion, $d\Xi_{p}$ is surjective, and $\mathscr{V}=\ker(d\Xi)$ is a smooth subbundle.
@@
    $d\Xi_{p}|_{\mathscr{H}_{p}}:\mathscr{H}_{p}\to T_{\Xi(p)}\mathcal{B}$ is a linear isomorphism.
@@
Consider a pseudo-Riemannian metric $g_{\mathcal{M}}$ on $\mathcal{M}$ such that $g_{\mathcal{M}}|_{\mathscr{V}_{p}}$ is non-degenerate. This metric induces an Ehresmann connection by defining
@@
where $\beta=1,\ldots,n+1$, for a basis $\{w_{\beta}\}$ of $\mathscr{V}_{p}$. The matrix $g_{\mathcal{M}}(w_{\alpha},w_{\beta})$ is invertible since $g_{\mathcal{M}}|_{\mathscr{V}_{p}}$ is non-degenerate, ensuring a unique $v$.
@@
$d\Xi_{p}\left(T_{p}\mathcal{M}\right)=d\Xi_{p}\left(\mathscr{H}_{p}\right).$ (47)
Given $\dim\mathscr{H}_{p}=\dim T_{\Xi(p)}\mathcal{B}=n$ and $\Xi$ a submersion, $d\Xi_{p}|_{\mathscr{H}_{p}}$ is an isomorphism. The smoothness of $\mathscr{H}$ follows from the smoothness of $g_{\mathcal{M}}$ and $\Xi$. ∎
@@
The construction of a metric with the previous properties will be based on the following considerations. Earlier, it was argued that each fiber $\Xi^{-1}(\sigma)$ is diffeomorphic to $\mathbb{R}^{n+1}$. We can further establish such fibers to be graphs of smooth functions over the variables $(S,\boldsymbol{a})$. We now develop the essential idea behind this fact: since the functions $\mu_{1},\dots,\mu_{n}$ satisfy the so-called joint injectivity condition, for a given fiber $\Xi^{-1}(\sigma)$ each point of the plane $(S,\boldsymbol{a})$ corresponds uniquely to a point of the fiber, given by the solution of the system of equations $\mu_{1}(S,\boldsymbol{a},\boldsymbol{\lambda})=c_{1},\ldots,\mu_{n}(S,\boldsymbol{a},\boldsymbol{\lambda})=c_{n}$. Such correspondence defines a map $\varphi_{\sigma}:\mathbb{R}^{n+1}\longrightarrow\Xi^{-1}(\sigma)$. Moreover, due to the submersion condition, the inverse function theorem guarantees that $\varphi_{\sigma}$ is a diffeomorphism, so that the fiber is realized as the smooth graph
$\Xi^{-1}(\sigma)=\left\{\left(S,\boldsymbol{a},\pi\circ\varphi_{\sigma}(S,\boldsymbol{a})\right)\;\big|\;S,\boldsymbol{a}\in\mathbb{R}^{n+1}\right\},$ (48)
where $\pi$ is the projection $(S,\boldsymbol{a},\boldsymbol{\lambda})\longmapsto\boldsymbol{\lambda}$. In particular, we can use the variables $(S,\boldsymbol{a})$ to furnish a coordinate chart of each fiber, allowing to define metrics and other tensor fields on the fiber in a simple and explicit manner.
In what follows, it will be useful to enlarge the previous construction to a full coordinate chart on the quantum thermodynamic fiber bundle. Given an element $\in\mathcal{M}, pp$ belongs to a unique fiber. Under the assumption that the map $\boldsymbol{\lambda}\mapsto\rho_{\boldsymbol{\lambda}}$ is injective, we can assign to $p$ the values $\boldsymbol{\lambda}$ of the only point where the fiber intersects the equilibrium submanifold $\mathcal{E}$. Let us call these $\bar{\boldsymbol{\lambda}}$. Then, we use the diffeomorphism $\varphi_{\sigma}$, where $\sigma$ here denotes the corresponding Gibbs state. Specifically, we compute $\varphi_{\sigma}^{-1}(p)\in\mathbb{R}^{n+1}$ and denote the result by $(\bar{S},\bar{\boldsymbol{a}})$. In this manner, each point $p\in\mathcal{M}$ is mapped to the $2n+1$ real numbers $(\bar{S},\bar{\boldsymbol{a}},\bar{\boldsymbol{\lambda}})$, and this mapping is a diffeomorphism by virtue of the preceding remarks. The new coordinates are adapted to the fiber bundle structure in the sense that two points $p,q\in\mathcal{M}$ belong to the same fiber if any only if their $\bar{\boldsymbol{\lambda}}$ coordinates coincide. We are now ready to provide the expression of a suitable metric on $\mathcal{M}$:
@@
A pseudo-Riemannian metric on $\mathcal{M}$, $g_{\mathcal{M}}$, satisfying the former desiderata is
$g_{\mathcal{M}}=g_{\bar{S}}d\bar{S}^{2}+\sum_{i=1}^{n}g_{\bar{a}_{i}}d\bar{a}_{i}^{2}+\Xi^{*}g_{\mathrm{BW}}+\sum_{i=1}^{n}h_{i}(d\bar{S}\otimes d\bar{\lambda}_{i}+d\bar{\lambda}_{i}\otimes d\bar{S}),$ (49)
where $g_{\bar{S}}\in\mathbb{R}$ controls the entropy direction, which may be negative; $g_{\bar{a}_{i}}>0$ are positive-definite on expectation value directions, for all $i=1,\ldots,n$; $\Xi^{*}g_{\mathrm{BW}}=\sum_{i,j}g_{ij}d\bar{\lambda}_{i}d\bar{\lambda}_{j}$ pulls back the Bures-Wasserstein metric; and finally $h_{i}$ are cross-terms ensuring contact compatibility, for all $i=1,\ldots,n$.
Note that, on $\mathcal{E}$
$g_{\mathcal{M}}|_{T\mathcal{E}}=\Xi^{*}g_{\mathrm{BW}},$ (50)
whereas on a fiber $F_{\sigma}$,
$g_{\mathcal{M}}|_{F_{\sigma}}=g_{\bar{S}}d\bar{S}^{2}+\sum_{i=1}^{n}g_{\bar{a}_{i}}d\bar{a}_{i}^{2}.$ (51)
For a path $\gamma:[0,T]\to\mathcal{M}$ with $\gamma(t)=(\bar{S}(t),\bar{\boldsymbol{a}}(t),\bar{\boldsymbol{\lambda}}(t))$, the thermodynamic length is
$L(\gamma)=\int_{0}^{T}\sqrt{|g_{\mathcal{M}}(\gamma^{\prime}(t),\gamma^{\prime}(t))|}dt,$ (52)
@@
$g_{\mathcal{M}}(\gamma^{\prime},\gamma^{\prime})=g_{\bar{S}}\bar{S}^{\prime 2}+\sum_{i=1}^{n}g_{\bar{a}_{i}}(\bar{a}_{i}^{\prime})^{2}+\sum_{i,j=1}^{n}g_{ij}\bar{\lambda}_{i}^{\prime}\bar{\lambda}_{j}^{\prime}+2\sum_{i=1}^{n}h_{i}\bar{S}^{\prime}\bar{\lambda}_{i}^{\prime}.$ (53)
Along a relaxation path $\gamma(t)\in F_{\sigma}$
$L(\gamma)=\int_{0}^{T}\sqrt{\left|\displaystyle g_{\bar{S}}\bar{S}^{\prime 2}+\sum_{i=1}^{n}g_{\bar{a}_{i}}(\bar{a}_{i}^{\prime})^{2}\right|}dt.$ (54)
@@
For a smooth path $\gamma:[0,1]\to\mathcal{B}$, with $\gamma(0)=\sigma_{1}$, $\gamma(1)=\sigma_{2}$, and $p_{0}\in\Xi^{-1}(\sigma_{1})$, the horizontal lift is a curve $\tilde{\gamma}:[0,1]\to\mathcal{M}$ satisfying